OpenStudy (anonymous):

cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (loser66):

hahaha. the problem was in 1 month ago!!!!??? still need help???

OpenStudy (zzr0ck3r):

how Is that funny?

OpenStudy (loser66):

To me, if I am stuck and need help, I come here to make question. If I don't get answer in 3 days, the problem is solved in class for sure. How can it last 1 month?

OpenStudy (anonymous):

I still need it to be explained because I'm going over the things I didn't understand.

OpenStudy (anonymous):

If someone could explain how to do this that would be wonderful

OpenStudy (loser66):

ok, I am so sorry friend, let me see whether I can help or not.

OpenStudy (loser66):

yours is \[cot xsec^4x = cotx -2 tan x +tan^3 x\] Prove or solve?

OpenStudy (anonymous):

Prove

OpenStudy (loser66):

\[sec^4x = (sec^2x)^2 = (1+tan^2x)^2\]

OpenStudy (loser66):

and then \[(1+tan^2x)^2 = 1 +2 tan^2x + tan^4 x\]

OpenStudy (loser66):

you still have cot x in the front of them, now, let the whole thing be \[cot x (1 +2 tan^2x +tan^4 x= cot x +2 cot x tan^2 x +cot x tan^4x\] on the other hand, you have \[cot x= \frac{1}{tanx}\] therefore yours = \[cotx +2 tanx +tan^3 x\] the problem is proven

OpenStudy (loser66):

I compensate for my fault when laughing at the time of the question. Forgive me?

OpenStudy (anonymous):

Thank you :)

OpenStudy (loser66):

Please, feel free to make question if there is any